OpenStudy (anonymous):

1. If f'(1/3*x^3)=2x^5, what is f'(x)? 2. Find the limit lim n->infinity (1/(2n+1)+1/(2n+2)+...+1/(2n+n) by computing a suitable definite integral. I am stressed out, thank you a lot for you help :) x

OpenStudy (turingtest):

for the first one, integrate 2x^5 to get f(1/3*x^3), from which f(x) should be obvious from there getting f'(x) is easy

OpenStudy (anonymous):

@TuringTest Not so fast! ;P \(\large\displaystyle \int f^{\prime}\left(\frac{x^3}{3}\right)\,dx \neq f\left(\frac{x^3}{3}\right)\). You need to take into consideration the \(\large \frac{1}{3}x^3\) term inside \(\large f^{\prime}\). Therefore, you have to make a substitution first before integrating.